What mass of silver nitrate AgNo3, is needed to make 0.50L of a 0.10mol/L solution?


1
Expert's answer
2021-05-07T05:25:57-0400

Let "x" be number of moles in 0.50L of solution , to find the value of "x" molarity of both the solution will be same "\ 0.10=\dfrac{x}{0.50}" ,"\ x=0.05\ mol"

Now you can calculate mass of silver nitrate , n = "\dfrac{m}{Mw}" , Where n = number of moles

m = mass

Mw = molecular weight = Molecular weight of "AgNO_3=169.87\ MW"

"m=n\times Mw=" 169.87"\times0.05=" 8.4935 g